State of Nature

Alternatives

A

B

C

Alternative 1

100

120

180

Alternative 2

200

100

50

Alternative 3

120

140

120

Do Nothing

0

0

0

The probabilities for states of nature A, B, and C are 0.3, 0.5, and 0.2, respectively. If a perfect forecast of the future were available, what is the expected value of perfect information (EVPI)?
Answer #1

expected value of alt 1 =(100*0.3+120*0.5+180*0.2)=126

expected value of alt 2 =(200*0.3+100*0.5+50*0.2)=120

expected value of alt 3 =(120*0.3+140*0.5+120*0.2)=130

therefore expected value without information =max(126,120,130) =130

expected value with perfect information =200*0.3+140*0.5+180*0.2=166

therefore expected value of perfect information (EVPI) =166-130 =36